# Jaike van Twiller1    Yossiri Adulyasak2    Erick Delage2    Djordje Grbic1    Rune Møller Jensen1

### v2

@@ 2 Domain Preliminaries
Utilization. Vessel utilization $u_{p}\in\mathbb{R}^{|B|\times|D|\times|K|\times|\textit{TR}|}$ represents the cargo placement across bays $B$, decks $D$, cargo types $K$, and transports TR. Load operations are defined by $u^{+}_{p}\in\mathbb{Z}_{>0}^{|B|\times|D|\times|K|\times|\textit{TR}|}$, whereas discharging is denoted by $u^{-}_{p}\in\mathbb{Z}_{>0}^{|B|\times|D|\times|K|\times|\textit{TR}|}$. The utilization at port $p$ is defined as $u_{p}=u_{p-1}+u^{+}_{p}-u^{-}_{p}\;\forall p\in{P}$, where $u_{0}$ is the vessel’s arrival condition at the first port. We also define the vessel’s pre-loading utilization $u^{\prime}_{p}=u_{p-1}-u^{-}_{p}$.
@@ 4.1 Formal MDP
Let ${\mathcal{M}}=({S},{X},\mathcal{T},{\mathcal{R}},{P}^{N_{P}-1}_{1},\gamma)$ define an episodic discounted MDP representing the MPP, where ${S}$ is a set of states, ${X}$ is a set of actions, $\mathcal{T}:{S}\times{X}\rightarrow\Delta({S})$ is the transition function, ${\mathcal{R}}:{S}\times{X}\times{P}^{N_{P}-1}_{1}\rightarrow\mathbb{R}$ is the reward function, ${P}^{N_{P}-1}_{1}$ is the finite horizon of load ports with $N_{P}-1$ being the last port, and $\gamma$ is the discounting factor.
State. The state is given by ${s}_{p}\in{S}$, defined as ${s}_{p}=(u_{p},q_{p},\zeta)$. This includes vessel utilization $u_{p}\in\mathbb{R}_{\geq 0}^{n_{u}}$ and realized demand $q_{p}\in\mathbb{R}_{\geq 0}^{n_{q}}$, where $n_{u}=|B|\times|D|\times|K|\times|\textit{TR}|$, $n_{c}=|B|\times|D|$ and $n_{q}=|K|\times|\textit{TR}|$ are the elements in utilization, location and demand, respectively.
@@
Transition. We use a stochastic transition function $\mathcal{T}({s}_{p+1}|{s}_{p},{x}_{p},)\in\Delta({S})$. Within an episode, the transition consists of multiple components:
@@ D.2 MIP Formulation
  • Hatch movement: $\tilde{hm}^{\phi}_{p,b}\in\{0,1\}$
In addition, we introduce the big $\mathbf{M}$ as a sufficiently large constant used to enforce logical constraints.
@@
$\sum_{k\in{K}}\sum_{\textit{tr}\in\textit{TR}^{M}_{p}}\tilde{u}^{b,d_{h},\phi}_{\textit{tr},k}\leq\mathbf{M}\tilde{hm}^{\phi}_{p,b}$
@@
$\sum_{k\in{K}}\sum_{\textit{tr}\in\textit{TR}^{\textit{AC}}(p)}\tilde{u}^{b,d_{o},\phi}_{\textit{tr},k}-M(1-\tilde{hm}^{\phi}_{p,b})\leq\tilde{\textit{ho}}^{\phi}_{p,b}$
@@
${\overline{z}}^{\phi}_{p}=(1+\delta^{\textit{cm}})\frac{2}{|{B}|}\sum_{\textit{tr}\in\textit{TR}^{M}_{p}}\sum_{k\in{K}}q^{\phi}_{\textit{tr},k}$
@@
${\overline{z}}^{\phi}_{p}-\sum_{b\in{b^{\prime}}}\sum_{d\in{D}}\sum_{k\in{K}}\sum_{\textit{tr}\in\textit{TR}^{M}_{p}}\tilde{u}^{b,d,\phi}_{\textit{tr},k}\leq\tilde{\textit{cm}}^{\phi}_{p}$
@@
$\qquad\qquad\qquad\forall p\in{P},\textit{tr}\in\textit{TR}^{\textit{L}}(p),k\in{K},$

### v3

@@ 2 Domain Preliminaries
Utilization. Vessel utilization $u_{p}\in\mathbb{Z}^{|B|\times|D|\times|K|\times|\textit{TR}|}$ represents the cargo placement across bays $B$, decks $D$, cargo types $K$, and transports TR. Load operations are defined by $u^{+}_{p}\in\mathbb{Z}_{>0}^{|B|\times|D|\times|K|\times|\textit{TR}|}$, whereas discharging is denoted by $u^{-}_{p}\in\mathbb{Z}_{>0}^{|B|\times|D|\times|K|\times|\textit{TR}|}$. The utilization at port $p$ is defined as $u_{p}=u_{p-1}+u^{+}_{p}-u^{-}_{p}\;\forall p\in{P}$, where $u_{0}$ is the vessel’s arrival condition at the first port. We also define the vessel’s pre-loading utilization $u^{\prime}_{p}=u_{p-1}-u^{-}_{p}$.
@@ 4.1 Formal MDP
Let ${\mathcal{M}}=({S},{X},\mathcal{T},{\mathcal{R}},{P}^{N_{P}-1}_{1},\gamma)$ define an episodic discounted MDP representing the MPP, where ${S}$ is a set of states, ${X}$ is a set of actions, $\mathcal{T}:{S}\times{X}\rightarrow\Delta({S})$ is the transition function, ${\mathcal{R}}:{S}\times{X}\times{P}^{N_{P}-1}_{1}\rightarrow\mathbb{R}$ is the reward function, ${P}^{N_{P}-1}_{1}$ is the finite horizon of load ports with $N_{P}-1$ being the last port, and $\gamma\in(0,1]$ is the discounting factor.
State. The state is given by ${s}_{p}\in{S}$, defined as ${s}_{p}=(u_{p},q_{p},\zeta)$. This includes vessel utilization $u_{p}\in\mathbb{R}_{\geq 0}^{n_{u}}$ and realized demand $q_{p}\in\mathbb{R}_{\geq 0}^{n_{q}}$, where $n_{u}=|B|\times|D|\times|K|\times|\textit{TR}|$, $n_{c}=|B|\times|D|$ and $n_{q}=|K|\times|\textit{TR}|$ are the shapes of utilization, location and demand, respectively.
@@
Transition. We use a stochastic transition function $\mathcal{T}({s}_{p+1}|{s}_{p},{x}_{p})\in\Delta({S})$. Within an episode, the transition consists of multiple components:
@@ D.2 MIP Formulation
  • Hatch movement: $\tilde{\textit{hm}}^{\phi}_{p,b}\in\{0,1\}$
In addition, we introduce the big $M$ as a sufficiently large constant used to enforce logical constraints.
@@
$\sum_{k\in{K}}\sum_{\textit{tr}\in\textit{TR}^{M}(p)}\tilde{u}^{b,d_{h},\phi}_{\textit{tr},k}\leq M\tilde{\textit{hm}}^{\phi}_{p,b}$
@@
$\sum_{k\in{K}}\sum_{\textit{tr}\in\textit{TR}^{\textit{AC}}(p)}\tilde{u}^{b,d_{o},\phi}_{\textit{tr},k}-M(1-\tilde{\textit{hm}}^{\phi}_{p,b})\leq\tilde{\textit{ho}}^{\phi}_{p,b}$
@@
${\overline{z}}^{\phi}_{p}=(1+\delta^{\textit{cm}})\frac{2}{|{B}|}\sum_{\textit{tr}\in\textit{TR}^{M}(p)}\sum_{k\in{K}}q^{\phi}_{\textit{tr},k}$
@@
${\overline{z}}^{\phi}_{p}-\sum_{b\in{b^{\prime}}}\sum_{d\in{D}}\sum_{k\in{K}}\sum_{\textit{tr}\in\textit{TR}^{M}(p)}\tilde{u}^{b,d,\phi}_{\textit{tr},k}\leq\tilde{\textit{cm}}^{\phi}_{p}$
@@
$\qquad\qquad\qquad\forall p\in{P},\textit{tr}\in\textit{TR}^{+}(p),k\in{K},$